# Wake Vision: A Tailored Dataset and Benchmark Suite for TinyML Computer Vision Applications

### v3

@@ Abstract
Tiny machine learning (TinyML) for low-power devices lacks robust datasets for development. We present Wake Vision, a large-scale dataset for person detection that contains over 6 million quality-filtered images. We provide two variants: Wake Vision (Large) and Wake Vision (Quality), leveraging the large variant for pretraining and knowledge distillation, while the higher-quality labels drive final model performance. The manually labeled validation and test sets reduce error rates from 7.8% to 2.2% compared to previous standards. In addition, we introduce five detailed benchmark sets to evaluate model performance in real-world scenarios, including varying lighting, camera distances, and demographic characteristics. Training with Wake Vision improves accuracy by 1.93% over existing datasets, demonstrating the importance of dataset quality for low-capacity models and dataset size for high-capacity models. The dataset, benchmarks, code, and models are available under the CC-BY 4.0 license, maintained by the Edge AI Foundation.
@@ 1 Introduction
Ultra-low power machine learning (TinyML) has emerged as an important technology to enable efficient ML deployments by co-locating models with sensors on microcontrollers (MCUs) [3, 38, 1, 6]. Although this approach dramatically reduces energy consumption and deployment costs, it imposes severe constraints. For instance, the models must often be within hundreds of kilobytes of memory in size, nearly four orders of magnitude less than mobile phones[2].
TinyML requires large-scale datasets to facilitate high-quality research. However, the extreme constraints imposed by the hardware necessitate that TinyML models be compact and efficient, rendering them unsuitable for traditional, complex ML tasks. As a result, conventional ML datasets like ImageNet [7], designed for more resource-intensive models, are often ill-suited for TinyML research endeavors. While existing TinyML datasets [5, 37] have been valuable, they tend to be limited in scale, making it challenging to train production-grade TinyML models. This creates a fundamental challenge where we can not develop robust TinyML models due to existing datasets being either too complex for tiny architectures or too limited in scale and diversity.
Table 1: A comparison of person detection and image classification datasets.
@@
Our work – Wake Vision (Quality) 1,322,574 624,115 9,291 27,881
Our work – Wake Vision (Large) 5,760,428 2,880,214 -- --
@@
To address this critical gap, we present Wake Vision, a large-scale dataset specifically designed for person detection, the quintessential vision use case for TinyML [1]. Person detection is a binary image classification task that, while simple, enables crucial applications ranging from occupancy detection [31] and smart HVAC/lighting [39] to acting as an always-on ’wake model’ in larger ML systems [15].
Wake Vision advances the state-of-the-art by providing 6M images, approximately 100x more than the current leading dataset Visual Wake Words (VWW) [5], while achieving a 1.93% accuracy improvement. The dataset is derived from Open Images v7 [21, 19] and is released under a permissive CC-BY 4.0 license to facilitate broad adoption.
A key contribution of our work is the first systematic investigation of how dataset quality versus quantity impacts TinyML model performance. This question is particularly relevant for resource-constrained models, as prior work suggests that large, overparameterized models can effectively identify and adapt to training set errors [40, 9]. TinyML models, with their limited capacity, may exhibit fundamentally different behavior. To explore this hypothesis, we introduce two complementary training sets within the Wake Vision visual corpus: Wake Vision (Large), emphasizing scale, and Wake Vision (Quality), prioritizing label accuracy. Our experiments reveal that while Wake Vision (Quality) outperforms its larger counterpart in naive training scenarios, optimal performance is achieved through a training strategy that leverages both sets — using Wake Vision (Large) for pre-training followed by using Wake Vision (Quality) fine-tuning. To further explore the dataset quality and quantity trade-off for tinyML, we train a range of ResNet [14] style models across Wake Vision datasets of varying error rates.
Our experiments collectively reveal that small models are more sensitive to higher error rates than large models and that large models can benefit from larger datasets despite a higher error rate. These findings provide insight into the unique challenges of training highly constrained models.
Beyond raw performance metrics, real-world deployment of TinyML systems often demands robustness across diverse and challenging use case conditions. Internet-sourced datasets often underrepresent challenging scenarios, such as low-lighting conditions, leading to overly optimistic performance estimates based on standard test set metrics.
To enable a rigorous evaluation of real-world capabilities, we additionally develop a comprehensive suite of five fine-grained benchmarks based on our Wake Vision dataset. Our benchmarks systematically evaluate performance across multiple dimensions: distance (detection at varying ranges), lighting (performance in well-lit and poorly-lit environments), depictions (handling of drawings and sculptures), perceived gender (investigating potential gender biases), and perceived age (examining effectiveness across age groups).
This benchmarking framework addresses a gap in current TinyML assessment methods and enables researchers to identify potential biases, limitations, and areas for improvement that might be masked by aggregate metrics alone. We find that Wake Vision models outperform VWW on 13 of the 16 fine-grained sets, demonstrating robust performance further illustrating the advantages of a large diverse training dataset.
In summary, we present the following contributions:
  • A production grade computer vision dataset for TinyML containing over 6M images (100$\times$ larger than state-of-the-art) for person detection, with comprehensive filtering, permissive licensing, and accessibility through standard ML frameworks.
  • A cleaned test set with 2.2% label error rate, enabling more reliable evaluation of TinyML models.
  • A benchmark suite for systematically evaluating model robustness across five critical dimensions (perceived gender, age, distance, lighting, depictions), exposing potential failure modes in real-world deployments.
  • Evidence that TinyML models are more sensitive to dataset quality than larger models, providing guidance for dataset curation in resource-constrained settings.
  • A complete, maintained dataset infrastructure including benchmarks, evaluation code, and reproducible filtering pipeline to support ongoing TinyML research.
@@
Figure 1: Sample images from the Wake Vision fine-grained benchmark suite.
@@ 2 Background and Related Work
Person Detection and Its Role in TinyML
Person detection is a fundamental computer vision task whose job is to determine the presence or absence of humans in visual data. Unlike more complex tasks such as pose estimation or instance segmentation, person detection focuses on the binary classification problem.
This seemingly simple task has emerged as a canonical use case for TinyML systems [3, 1], driven by both theoretical foundations and practical considerations. In the context of resource-constrained computing, TinyML systems demand tasks that carefully balance computational feasibility with real-world utility. Person detection represents an optimal compromise. It is sufficiently lightweight to be implemented on embedded microcontrollers while maintaining enough discriminative power for practical applications.
The significance of person detection extends beyond its computational characteristics, serving as a critical selective activation mechanism in resource-constrained systems. These lightweight models efficiently manage the activation of power-intensive sensors and processors by determining human presence. This architectural pattern has demonstrated particular value in diverse applications, including occupancy detection [31], smart HVAC systems [39], and as an always-on "wake model" in larger ML pipelines [15]. Moreover, this approach enables substantial energy savings while preserving privacy through on-device processing.
@@
In the TinyML computer vision domain, the VWW dataset [5, 6, 24, 2, 1] has established itself as the de facto standard. It remains the only open-source dataset specifically designed for person detection with direct commercial applicability and potential for real-world impact. However, VWW faces significant limitations: its small size and indirect access requirements (requiring regeneration from MS-COCO [25]) restrict its accessibility and utility.
@@
3 Wake Vision: Dataset and Benchmark Suite
We introduce Wake Vision, a new dataset for TinyML person detection. The dataset’s size, quality, and detailed metadata enable new avenues of TinyML research. Wake Vision is almost two orders of magnitude larger than prior comparable datasets and can be adapted based on the target use case. Fig. 1 illustrates examples of images in the dataset. In this section, we discuss our methodologies for label generation, data filtering, and error correction, as well as discussing the usability of the data set and comparing it with VWW Dataset.
@@ 3.1 Label Generation
A large person detection dataset is indispensable for TinyML research. It can, however, be challenging to obtain. Manually labeling millions of images would be prohibitively expensive for a nascent field like TinyML. Therefore, we bootstrap Wake Vision using existing large-scale data efforts.
The base label in Wake Vision is a binary person/non-person label. We derive these labels from Open Images [21, 19], which contain both image-level and bounding box labels of various objects. Image-level labels describe which objects are present in an image but not localized to a specific point in the image. Bounding box labels provide a set of four coordinates forming a box around the object.
@@
The bounding box label classes are hierarchically structured so that one class can be a subcategory or a part of another class. For example a “Woman” is a subcategory of a “Person,” and a “Human Hand” is a part of “Person.” A comprehensive list of Open Images labels related to Wake Vision can be found in Appendix L.
@@
Users can adapt Wake Vision to meet the needs of specific use cases by adding several configurable options to the open-source dataset creation script. For example, a configuration option exists to change whether depictions are labeled as non-persons or excluded from the dataset. We refer to Appendix I for details about these configuration options.
@@
We observe that the bounding box labels are generally more accurate and provide more information for data filtering but are less numerous. The Open Images training set has ~9 million images with image-level person labels, but only ~1.7 million images with a person bounding box. Consequently, this presents a trade-off between more data (image-level labels) or higher-quality labels (bounding box labels). Because of this trade-off, we provide two Wake Vision training sets: Wake Vision (Large) and Wake Vision (Quality). One larger set, labeled via Open Images image-level labels, and one smaller set, labeled via Open Images bounding box labels (Table 1). The Open Images validation and test sets are fully labeled with bounding box labels. Therefore, both the Wake Vision validation and test sets are created using the higher-quality bounding box labels.
Large Training Set from Image-Level Labels
@@
Image-level labels in Open Images contain little additional information than the raw labels. The only additional relevant information is a confidence property, which represents how certain it is that a label is correct. The confidence ranges from 0 to 1 for labels generated by machines and is either 0 or 1 for labels negatively or positively verified by humans, respectively. See Sec. 3.2 for more information. We use this property to exclude images from Wake Vision if they contain only low-confidence person labels.
Quality Training Set From Bounding Box Labels
@@
Bounding Box labels in Open Images, in contrast to image-level labels, are all verified and localized by humans. This ensures that bounding box labels are less likely to be false positives. Bounding box labels can be used to calculate an approximation of the area taken up by a person. We use this as a proxy for the distance of a person in the image, and by default exclude images where all person labels are far away, i.e., take up less than 5% of the image. An example of an image with a faraway person can be found in Fig. 2(a). Finally, bounding box attributes include a depiction flag, which represents whether a label is a depiction. See an example of an image containing a person with the depiction flag set in Fig. 2(b). By default, we use this flag to make sure that we do not consider a depiction a person. See Appendix C for an illustration of the bounding box filtering process.
@@ 3.2 Label Correction
Label errors are a challenge in computer vision that limit progress and obscure true performance [30, 4]. This is especially prevalent in large datasets that use complex labeling pipelines to scale while managing costs. Recognizing the importance of label quality, we take measures to estimate and improve the accuracy of its labels, aligning with best practices in the field of computer vision.
Wake Vision’s labels are derived from the Open Images labels; therefore, the errors that occur in the Open Images labeling pipeline are inherited by Wake Vision. All labels in the Open Images dataset originate from machine-generated candidate labels. Many of these labels are later verified by human annotators, and a subset of them are given bounding boxes and additional annotations.
While the machine-generated label phase aims to identify objects efficiently, any instances missed during this initial step are unlikely to be captured in downstream phases. Consequently, the Open Images dataset may contain numerous false negative labels, referring to images where an object is present but lacks the corresponding label. Therefore, additional measures are necessary to identify and correct such labeling omissions to ensure the dataset’s integrity.
Given Wake Vision’s size, manually correcting label errors across the entire dataset becomes an arduous undertaking. Therefore, we prioritize labeling the Wake Vision validation and test sets so that the dataset can be used to accurately evaluate a model’s performance.
We use the Scale Rapid tool from Scale AI (https://scale.com/) to relabel the Wake Vision validation and test sets. Details about the relabelling process, including instructions given and total cost, can be found in Appendix N.
With the ground truth established in the validation and test sets, Wake Vision becomes an ideal foundation for research on automated data cleaning techniques [29]. We perform an initial exploration of these methods in appendix O.
We report our estimated label error rate after label corrections versus VWW in Table 2. These estimates are all based on a random subset of 500 samples from each dataset, which the authors manually relabelled according to whether we saw a non-depiction person or body part in the image. The Wake Vision validation and test set is considerably lower than that of the VWW dataset top-level error rate.
Table 2: Label error rate of VWW and Wake Vision after label correction. The Wake Vision error rate is broken down by set. The Wake Vision Train (Quality) and Train (Large) have different error rates due to the different label sources. The Wake Vision Validation and Test sets have a lower error rate due to manual relabeling.

### v5

@@ Abstract
Tiny machine learning (TinyML) for low-power devices lacks systematic methodologies for creating large, high-quality datasets suitable for production-grade systems. We present a novel automated pipeline for generating binary classification datasets that addresses this critical gap through several algorithmic innovations: intelligent multi-source label fusion, confidence-aware filtering, automated label correction, and systematic fine-grained benchmark generation. Crucially, automation is not merely convenient but necessary to cope with TinyML’s diverse applications. TinyML requires bespoke datasets tailored to specific deployment constraints and use cases, making manual approaches prohibitively expensive and impractical for widespread adoption. Using our pipeline, we create Wake Vision, a large-scale binary classification dataset of almost 6 million images that demonstrates our methodology through person detection—the canonical vision task for TinyML. Wake Vision achieves up to a 6.6% accuracy improvement over existing datasets via a carefully designed two-stage training strategy and provides 100× more images. We demonstrate our broad applicability for automated large-scale TinyML dataset generation across two additional target categories, and show our label error rates are substantially lower than prior work. Our comprehensive fine-grained benchmark suite evaluates model robustness across five critical dimensions, revealing failure modes masked by aggregate metrics. To ensure continuous improvement, we establish ongoing community engagement through competitions hosted by the Edge AI Foundation. All datasets, benchmarks, and code are available under CC-BY 4.0 license, providing a systematic foundation for advancing TinyML research.
**footnotetext: These authors contributed equally to this work.
@@ 1 Introduction
Ultra-low power machine learning (TinyML) has emerged as an important technology to enable efficient ML deployments by co-locating models with sensors on microcontrollers (MCUs) [3, 47, 1, 8]. Although this approach dramatically reduces energy consumption and deployment costs, it imposes severe constraints that fundamentally differ from standard machine learning scenarios. For instance, the models must often fit within hundreds of kilobytes of memory, nearly four orders of magnitude less than common model sizes on smartphones[2]. These extreme hardware-imposed constraints necessitate compact and efficient TinyML models, rendering traditional Machine Learning (ML) benchmarks and datasets unsuitable. For example, ImageNet [10] models must output a thousand classes, which is infeasible on MCUs and even smartphone models struggle to support [7].
Existing TinyML datasets such as Visual Wake Words (VWW) [7] and Google Speech Commands [46] target simpler applications, requiring support for binary classification and tens of classes, respectively. While these datasets are valuable, their limited scale and label error rates make them unsuitable for pushing the frontier of TinyML research and training production-grade TinyML models. This creates a challenge where the development of robust TinyML systems is hindered by existing datasets being either too complex for tiny architectures or too limited in scale and diversity. However, this challenge extends beyond simply collecting more data—it requires systematic methodologies for rapidly creating, curating, and evaluating datasets that meet TinyML’s unique constraints at scale.
Table 1: A comparison of Wake Vision for person detection and image classification datasets.
@@
Ours – WV (Quality) 1,322,574 624,115 9,291 27,881
Ours – WV (Large) 5,760,428 2,880,214 -- --
@@
To address this gap, we present an automated dataset generation pipeline for creating and benchmarking large-scale binary classification datasets suitable for TinyML that incorporates several algorithmic innovations: intelligent fusion of multiple label sources, confidence-aware filtering strategies, automated label correction techniques, and systematic generation of fine-grained benchmark suites. TinyML dataset automation is not merely a convenience but a necessity—manual labeling approaches are prohibitively expensive (our manual validation of 70K images cost \$7,000, which would extrapolate to \$600,000 for a 6M image dataset) and suffer from inconsistency across labelers. Our dataset generation pipeline derives datasets from Open Images v7 [29, 27], which enables created datasets to be licensed under a permissive CC-BY 4.0 license to facilitate broad adoption.
We demonstrate the effectiveness of our pipeline by creating Wake Vision, a large-scale open-source binary classification dataset that uses person detection to validate our methodology. Person detection serves as the quintessential vision use case for TinyML [1] and is an ideal testbed because it enables several important applications ranging from occupancy detection [39] and smart HVAC/lighting [48] to acting as an always-on ‘wake model’ in larger ML systems [23]. As shown in Table 1, Wake Vision advances the state-of-the-art TinyML datasets by providing almost 6M images, close to 100$\times$ more than the current leading dataset VWW [7], while achieving up to a 6.6% accuracy improvement.
We use Wake Vision to derive insights about TinyML-specific design principles: small models benefit more from smaller, high-quality training sets compared to noisier, larger sets, but if designed carefully, two-stage training (pre-training on Wake Vision Large followed by fine-tuning on Wake Vision Quality) can leverage both scale and quality to achieve optimal performance (Section 4.1). Furthermore, to enable rigorous evaluation of real-world capabilities, we develop a comprehensive suite of five fine-grained benchmarks that systematically evaluate performance across critical dimensions including perceived gender and age, distance, lighting, and depictions (Section 3.3). This benchmarking framework addresses a gap in current TinyML assessment methods, revealing that Wake Vision trained models outperform VWW trained models on 13 of 16 fine-grained sets, demonstrating robust performance beyond aggregate metrics alone.
Beyond person detection, we show that our dataset generation pipeline generalizes to all 9.6K trainable classes provided by Open Images v7, enabling the systematic creation of more production-grade datasets across diverse TinyML applications (Section 3.5). We provide evidence with two additional domains for TinyML deployments: bird detection for smart nests [9] and automated feeders [20] (12$\times$ more images with label error rates improving from 6.4% to 4.8%), and car detection for parking monitoring [6] and blind-spot systems [43] (27$\times$ more images with error rates improving from 6.6% to 0.6%). Additionally, we host public challenges for long-term, community-driven enhancements to Wake Vision. Our dataset, benchmarks, code, and models, along with past and upcoming challenges, can be found at https://wakevision.ai.
@@
Figure 1: Sample images from the Wake Vision fine-grained benchmark suite (for person detection).
@@ 2 Background and Related Work
Generating Binary Classification Datasets for TinyML
Automatically generated datasets can significantly lower costs associated with deploying TinyML systems. Prior to our work, the primary way to generate binary classification datasets for TinyML applications was through the code used to create VWW [7], which can be applied to other MS-COCO [33] classes. While we offer the same capability for Open Images, Wake Vision goes beyond dataset generation, offering customizable curation properties comprising of label correction, fine-grained benchmark suite generation, and community-driven enhancements through competitions, addressing the whole dataset life-cycle.
Person Detection in TinyML Applications
Person detection has emerged as the canonical computer vision use case for TinyML systems [3, 1]. In the context of resource-constrained computing, TinyML systems demand tasks that carefully balance computational feasibility with real-world utility. Person detection represents an optimal compromise. It is sufficiently lightweight to be implemented on embedded microcontrollers while maintaining enough discriminative power for practical applications. In particular, person detection serves as a critical activation mechanism in resource-constrained systems. These lightweight models support always-on deployment on low-powered MCUs to selectively “wake” power-intensive sensors, processors, and larger, more capable ML models upon detecting human presence [23]. This architectural pattern enables substantial energy savings while preserving privacy through on-device processing, and supports diverse applications ranging from occupancy detection [39], smart HVAC systems [48], or poacher detection [11].
@@
In the TinyML computer vision domain, the VWW dataset [7, 8, 32, 2, 1] has established itself as the de facto standard. Before Wake Vision, it was the only open-source dataset specifically designed for person detection with direct commercial applicability. However, VWW faces significant limitations: its small size and indirect access requirements (requiring regeneration from MS-COCO [33]) restrict its accessibility and utility.
@@
3 Dataset and Benchmark Generation
With Wake Vision, we introduce a new dataset for TinyML person detection which is two orders of magnitude larger than prior datasets. The dataset’s size, quality, and detailed metadata enable new avenues of TinyML research.  Figure 1 illustrates example images in our dataset. This section presents our methodologies for dataset generation and curation, with a focus on person detection, including label generation, data filtering, error correction, and community-driven enhancements, as well as discussing the usability and generalization capabilities of our dataset pipeline in comparison with VWW.
@@ 3.1 Label Generation
A large person detection dataset is indispensable for TinyML research. However, manually labeling millions of images would be prohibitively expensive for a nascent field like TinyML. Therefore, we bootstrap Wake Vision using existing large-scale data efforts.
@@
The base label in Wake Vision is a binary person/non-person label. We derive these labels from Open Images [29, 27], which contain both image-level and bounding box labels of 9.6K trainable classes and 600 objects, respectively. Image-level labels are unlocalized, describing objects present in an image while bounding box labels localize an object by four coordinates. The bounding box label classes are hierarchically structured so that one class can be a subcategory or a part of another class. For example a “Woman” is a subcategory of a “Person,” and a “Human Hand” is a part of “Person.” A comprehensive list of Open Images labels related to Wake Vision can be found in Appendix L.
@@
Users can adapt Wake Vision to meet the needs of specific use cases using the many configuration options available in our binary classification dataset generation pipeline. For example, a configuration option exists to change whether artistic depictions of humans are labeled as non-persons or excluded from the dataset. We refer to Appendix I for details about these configuration options.
@@
We observe that bounding box labels are generally more accurate and provide more information for data filtering while being less numerous. The Open Images training set has ~9 million images with image-level person labels, but only ~1.7 million images with a person bounding box. Consequently, this presents a trade-off between more data (image-level labels) or higher-quality labels (bounding box labels). In response, we provide two Wake Vision training sets: Wake Vision (Large), labeled via Open Images image-level labels, and Wake Vision (Quality), a smaller set labeled via Open Images bounding boxes (Table 1). Since the Open Images validation and test sets are fully labeled with higher-quality bounding box labels, we derive the Wake Vision validation and test sets from the bounding box labels in these Open Images splits.
Wake Vision (Large) Training Set
@@
Image-level labels in Open Images include a confidence property, which represents how certain it is that a label is correct. This confidence ranges from 0 to 1 for machine-generated labels and is strictly either 0 or 1 for human-verified labels. See Section 3.2 for more information. We exclude low-confidence machine-labeled person images from Wake Vision.
Wake Vision (Quality) Training Set
@@
Bounding box labels in Open Images, in contrast to image-level labels, are all verified and localized by humans, minimizing false positive labels. Bounding box labels can be used to calculate an approximation of the image area taken up by a person. We use area as a proxy for the distance of a person to the camera, and by default exclude far away persons, i.e., take up less than 5% of the image (see  Figure 2(a) for an example). Finally, bounding box attributes include an artistic depiction flag (an example is provided in  Figure 2(b)). By default, Wake Vision does not consider a depiction a person. See Appendix D for an illustration of the bounding box filtering process.
@@ 3.2 Label Correction
Label errors are a challenge in computer vision that limit progress and obscure true performance [38, 5]. This is especially prevalent in large datasets that use complex labeling pipelines to scale while managing costs. Recognizing the importance of label quality, we take measures to estimate and improve the accuracy of Wake Vision’s labels.
As Wake Vision’s labels are derived from Open Images v7; errors in these labels are inherited by Wake Vision. All image-level labels in the Open Images dataset originate from machine-generated candidate labels. Many of these labels are later verified by human annotators, and a subset of them are given bounding boxes and additional annotations.
While the machine-generated label phase aims to identify objects efficiently, any instances missed during this initial step are unlikely to be captured in downstream phases. Consequently, the Open Images dataset may contain numerous false negative labels, referring to images where an object is present but lacks the corresponding label. Therefore, additional measures are necessary to identify and correct such labeling omissions. Given Wake Vision’s size, manually correcting label errors across the entire dataset becomes an arduous undertaking.
Therefore, we prioritize labeling the Wake Vision validation and test sets so that the dataset can be used to accurately evaluate a model’s performance. We use the Scale Rapid tool from Scale AI (https://scale.com/) to relabel the Wake Vision validation and test sets. Details about the relabelling process, including instructions given and total cost, can be found in Appendix N. With the ground truth established in the validation and test sets, Wake Vision becomes an ideal foundation for research on automated data cleaning techniques [37]. We perform an initial exploration of these methods in Appendix O.
We report our estimated label error rate after label corrections versus VWW in Table 2. These estimates are all based on a random subset of 500 samples from each dataset, manually evaluated by the authors. The Wake Vision validation and test set is considerably lower than that of the VWW dataset top-level error rate.
Table 2: Label error rate of VWW and Wake Vision after label correction. The Wake Vision Train (Quality) and Train (Large) have different error rates due to the different label sources. The Wake Vision Validation and Test sets have a lower error rate due to manual relabeling.